Two positive integers a and b are such that a+b = a/b + b/a. What is the value of a2 +b2

Grade:9

2 Answers

Arun
6 years ago
this is only valid if a = 1 and b = 1hence put a
Sriyansh Dash
5 years ago
a+b=a/b+b/a
=a+b=a^2+b^2/ab
=a^2b+b^2a=a^2+b^2
Comparing both sides we get ,
a^2b=a^2
=b=1
 
b^2a=b^2
=a=1